sec3x = sqrt2
solve for [0,2pi]

Answer by lwsshak3(11628)

sec3x = sqrt2
solve for [0,2pi]
sec3x = √2=1/cos3x
cos3x=1/√2=√2/2
3x=π/4, 5π/4
x=π/12, 5π/12